OpenStudy (anonymous):

Solve for Y=Ae^kt if Y=15.25 t=25 and Y=6.75 t=10 Find the value of A and K

OpenStudy (ivycoveredwalls):

Substitute values given for Y and t to get two equations: \[15.25=Ae ^{25k}\] \[6.75=Ae^{10k}\] Hint: Now divide the first equation by the second equation: \[\frac{15.25}{6.75}=\frac{Ae^{25k}}{Ae^{10k}}\] Can you use exponent properties to solve for k? Once you know the value of k, you can substitute it back in either of the first two equations to find A.

OpenStudy (anonymous):

what kind of problem are you doing

OpenStudy (anonymous):

@ivycoveredwalls is it 2.22 = e^3K ?

OpenStudy (michele_laino):

I think that: \[e ^{15k}=\frac{ 15.25 }{ 6.75 }\] as you easily can check using the equation that @ivycoveredwalls has wrote

OpenStudy (michele_laino):

so going to logarithm of the above expression, we can write: \[15*k=\ln(\frac{ 12.25 }{ 6.75 })\] and finally: \[k=\frac{ 1 }{ 15 }\ln(\frac{ 12.25 }{ 6.75 })\]

OpenStudy (anonymous):

is the answer K equal to 1.7039129684 ?

OpenStudy (michele_laino):

Sorry, I don't think, please apply the above formula!

OpenStudy (anonymous):

but why it is 12.25 above

OpenStudy (michele_laino):

because using the properties of the product of powers of the same base, we can write: \[\frac{ 12.25 }{ 6.75 }=\frac{ e ^{25k} }{ e ^{10k} }=e ^{25k-10k}=e ^{15k}\]

OpenStudy (anonymous):

I think it is 15.25 instead of 12.25 ?

OpenStudy (michele_laino):

Sorry, you are right!

OpenStudy (anonymous):

then the K is 0.0594665359?

OpenStudy (michele_laino):

@Kyle` that's right!

OpenStudy (michele_laino):

Sorry I got k=0.0543

OpenStudy (anonymous):

which one is correct :p?

OpenStudy (michele_laino):

I think k=0.0543

OpenStudy (anonymous):

how about A !?

OpenStudy (michele_laino):

In order to find A, we can use, for example, the second equation which @ivycoveredwalls has wrote, namely: \[6.75=A*e ^{10k}\] Now, we can write, using the properties of powers with the same base: \[e ^{10k}=(e ^{15k})^{2/3}\] because: \[15k*\frac{ 2 }{ 3 }=10k\] Since: \[e ^{15k}=\frac{ 15.25 }{ 6.75 }\] then, we can write: \[e ^{10k}=(\frac{ 15.25 }{ 6.75 })^{2/3}=1.72177\] finally we get for A the subsequent value: \[A=\frac{ 6.75 }{ 1.72 }\approx 3.92\]

OpenStudy (anonymous):

what is the 2/3 ?

OpenStudy (michele_laino):

2/3=0.666666....

OpenStudy (anonymous):

where is the 0.6666 come from !?

OpenStudy (michele_laino):

because to use the formula: 6.75=A*e^(10k), you have to calculate what is e^(10k). Now, we have e^10k), nothe that being: \[10k=\frac{ 2 }{ 3 }*15k\] we can get e^(10k) simply performing the subsequent calculus: \[e ^{10k}=(e ^{15k})^{(2/3)}\]

OpenStudy (michele_laino):

since, applying the rule of the power of power, we can write: \[(e ^{15k})^{(2/3)}=e ^{15k*\frac{ 2 }{ 3 }}=e ^{10k}\]

OpenStudy (anonymous):

thank you very much, its clear